OpenStudy (anonymous):

1000meters = 1 kilometer, 1/1000 meters =?

OpenStudy (anonymous):

1 millimeter

jimthompson5910 (jim_thompson5910):

1000meters = 1 kilometer Divide both sides by 1000 to get 1000/1000 meters = 1/1000 kilometer 1 meter = 1/1000 kilometer Divide both sides by 1000 again to get 1/1000 meters = (1/1000)/1000 kilometer 1/1000 meters = 1/1000000 kilometer

jimthompson5910 (jim_thompson5910):

You can go even further and say that 1/1000000 km = 1 mm So 1/1000 meters = 1 millimeter

OpenStudy (anonymous):

1/1000= 10^-3 m = 1 mm